OpenStudy (anonymous):

if a number is subtracted from 3, the principal square root of this difference is equal to the number decreased by 1 find the numbers?

Parth (parthkohli):

Ah, well, another word problem which translates to:\[\sqrt{3 - n} = n - 1 \]

Parth (parthkohli):

OK, so you square both sides... and then I need not tell the rest, right?

OpenStudy (anonymous):

yes i got it from here thanks

Parth (parthkohli):

You're welcome!